OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y =8/x^ + 4

OpenStudy (amistre64):

not real sure about the notation

OpenStudy (amistre64):

sqrt(..) helps to define the radical

OpenStudy (anonymous):

sorry, there shouldn't be a sqrt...

OpenStudy (amistre64):

a little better, what does /x^ +4 mean?

OpenStudy (anonymous):

8 divided by x squared plus 4

OpenStudy (amistre64):

\[1)~\frac{8}{x^2+4}\]\[2)~\frac{8}{x^2}+4\]

OpenStudy (anonymous):

2 opition :)

OpenStudy (amistre64):

now its clear :)

OpenStudy (amistre64):

\[f(x)=x+4~;~g(x)=\frac{8}{x^2}\] \[f(x)=\frac{8}{x}+4~;~g(x)={x^2}\] \[f(x)=\frac{8}{x^2}+4~;~g(x)=x\] is there anymore info to go on? otherwise you can create quite a few options

OpenStudy (amistre64):

\[f(x)=2(x+2)~;~g(x)=\frac{4}{x^2}\]

OpenStudy (anonymous):

one might imagine infinitely many

OpenStudy (anonymous):

Yiu should be able to have a few different choices, but none of those distribute an x to the 4?

OpenStudy (amistre64):

why would you need to distribute an x to the 4 ?

OpenStudy (anonymous):

You don't I'm just verifiyng.

OpenStudy (anonymous):

I see now! Sorry! Thank you so much :)

OpenStudy (amistre64):

any setup that combines g into the x part of f should suffice